# Multidimensional constructs and moderated linear and nonlinear factor analysis

### v1

@@ Discussion
Another area for development is the design of penalty functions themselves. While we focus here on ridge, lasso, and alignment penalties, other formulations may offer distinct advantages. Elastic net penalties, for example, combine the strengths of ridge and lasso by balancing smooth shrinkage with sparsity induction. Adaptive penalties can weight moderation effects differently, improving bias-variance tradeoffs. Group penalties could be used to enforce structure across related sets of moderation effects (e.g., all loadings associated with a particular factor). Systematic simulation studies will be essential to compare the performance of these alternatives under varying conditions of dimensionality, sample size, and degree of true non-invariance.

### v2

@@ Discussion
Another area for development is the design of penalty functions themselves. While I focus here on ridge, lasso, and alignment penalties, other formulations may offer distinct advantages. Elastic net penalties, for example, combine the strengths of ridge and lasso by balancing smooth shrinkage with sparsity induction. Adaptive penalties can weight moderation effects differently, improving bias-variance tradeoffs. Group penalties could be used to enforce structure across related sets of moderation effects (e.g., all loadings associated with a particular factor). Systematic simulation studies will be essential to compare the performance of these alternatives under varying conditions of dimensionality, sample size, and degree of true non-invariance.